OpenStudy (anonymous):

Solve the following equations. x^2+4=0 x^3+4x^2-21x=0 x^4-16=0 options 3,7,1 -7,0,3 2i,-2i 16,-16 2,-2,2i,-2i

hartnn (hartnn):

x^2+4=0 so, x^2 = -4 x^3+4x^2-21x=0 x (x^2) + 4 (x^2) -21x = 0 x(-4) + 4 (-4) -21x = 0 this is a linear equation in x :) can you solve it ?

OpenStudy (anonymous):

so I would substitute the x for -4? and then solve?

hartnn (hartnn):

x^2+4 =0 x^2 = -4 so you substitute x^2 as -4 as i did and you would solve for x from x(-4) + 4 (-4) -21x = 0

OpenStudy (anonymous):

so -4x-16-21x

OpenStudy (anonymous):

sorry I get confused easily..

hartnn (hartnn):

no problem :) yes thats correct so, -25 x = 16 but i don't see that in any choices :P

OpenStudy (anonymous):

so it would have to be 16,-16 wouldn't it?

hartnn (hartnn):

nopes, x^2 +4 = 0 gives you x^2 =- 4 x = \(\pm 2i\)

hartnn (hartnn):

x^4-16 =0 x^4 - 2^4 = (x^2+2^2)(x^2-2^2) = 0 so that x^2-2^2 = 0 gives you 2 other roots x^2 = 2^2 x= \(\pm 2 \)

OpenStudy (anonymous):

so 2,-2,2i,-2i..... I can't figure out the i equations

hartnn (hartnn):

yes thats correct you mean you didn't get how we get +2i and -2i ?

OpenStudy (anonymous):

I get it but I usually don't

hartnn (hartnn):

you can ask any doubts :)

OpenStudy (anonymous):

i will! will you still help me figure out the other two please? :) I think I get it?

hartnn (hartnn):

if i am online , i'll help :)

OpenStudy (anonymous):

okay so for x^3+4x^2-21x=0 I would enter in a option to see if it works right?

hartnn (hartnn):

if you have choices, than you can but easiest thing to do is to factor out 'x'

hartnn (hartnn):

x(x^2+4x-21) = 0 so, x=0 or x^2+4x-21 =0 so consider only those choices which have x=0

OpenStudy (anonymous):

they both equal 0

hartnn (hartnn):

so go for x^2+4x -21 =0 can you tell me 2 numbers whose sum is +4 and product is -21 ?

OpenStudy (anonymous):

-25 does

hartnn (hartnn):

?

OpenStudy (anonymous):

-25+4=-21

OpenStudy (anonymous):

isn't that what you asked?